If k × velocity × mass is displacement, then the dimensions of k are...
1
Expert's answer
2020-11-02T09:27:00-0500

Velosity has the dimensions "m\/s", mass - "kg", displacement - "m". let's denote the dimensions of "k" as "x". Then obtain the following equation:


"x\cdot \dfrac{m}{s}\cdot kg = m"

Expressing "x". obtian:


"x = m\cdot \dfrac{s}{m}\cdot \dfrac{1}{kg} = \dfrac{s}{kg}"

Answer. s/kg.